OpenStudy (anonymous):

Add and answer in simplest fractional form. = A. -11/24 B. 3/5 C. 3/24 D. 3/11

OpenStudy (anonymous):

What are the original fractions?

OpenStudy (anonymous):

im gonna show you hold on

OpenStudy (anonymous):

\[-\left(\begin{matrix}4 \\ 3\end{matrix}\right)\left(\begin{matrix}7 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

Are you sure? That gives an answer of \[-\frac{ 28 }{ 9 }\]

OpenStudy (anonymous):

it has an addition sign in the middle of the fractions sorry

OpenStudy (anonymous):

Well that just gives an answer of 1 :P

OpenStudy (anonymous):

-4/3 + 7/3 = 3/3 = 1.

OpenStudy (anonymous):

this is very confusing lol